The length of the chord of the parabola y2=x which is bisected at the point (2,1) is?


a)51/2     b)4*31/2     c)3*31/2     d)2*51/2

Grade:12

1 Answers

Jit Mitra
12 years ago

T = S1

y-0.5(x+2) = 1^2 - 2 = -1

 

2y - x - 2 = -2

or, x = 2y

 

Put in the equation,

 

y^2 = 2y

or, y = 0 , 2

x = 0, 4

Points where the line cuts the parabola = (0,0), (4,2)

 

Length = sqrt(4^2 + 2^2) = sqrt(16+4) = sqrt(20) = 2*sqrt(5)

 

Option (d)